CRITICAL CARE PROBLEMS

#10. Nesiritide is ordered to infuse at 0.01 mcg/kg/min for a patient with congestive heart failure. The nesiritide is available in a 1.5 mg/250mL D5W concentration. How many mL/hr should the nesiritide infuse for an 80-kg patient?

#11. A patient with supraventricular tachycardia has a maintenance infusion of Brevibloc 100 mcg/kg/min ordered. The Brevibloc concentration is 10mg/mL. The patient's weight is 80kg. How many mL/hr should the IV pump be set for?

#12. A patient with septic shock has a Neo-Synephrine drip infusing at 20mL/hr. The Neo-Synephrine concentration is 50mg in 250 mL of NS. How many mcg/min is the patient receiving?

#13. A patient in shock has an order for a Neo-Synephrine drip at 50 mcg/min. The pharmacy dilutes 50 mg of Neo-Synephrine in 250 mL of NS. How many mL/hr should the Neo-Synephrine drip be infused?

#14. Isuprel is ordered to infuse at 2 mcg/min for a patient with hypotension. The Isuprel is delivered to the unit in a 1 mg/250mL concentration. How many mL/hr should the Isuprel infuse?

#15. A patient with an allergy to heparin has an order for Refludan 0.1 mg/kg/h. The pharmacy dilutes 100mg of Refludan in 250mL of D5W. For a patient weighing 70kg, how many mL/hr should the Refludan drip be infused?
Answer #1

10.Given

1kg=0.01mcg of Nesiritide

So 80kg=??mcg

Cross multiply and divide

=80×0.01÷1

=0.8mcg/min

Available 1.5mg/250ml (150 mg=1500mcg)

So 0.8mcg=??ml

Cross multiply and divide

=250×0.8÷1500

=0.1 ml for 1 minute ,so for 1 hour multiply by 60

=6ml/hr

11.Weight =80 kg

Order:100mcg/kg/day

Therefore 80 kg=8000mcg /day(80×100)

Available 10mg/ml (1000mcg/ml)

There 8000mcg=??ml

Cross multiply and divide

=8000×1÷1000

=8ml/ min

So for 1 hour multiply by 60 minutes

=8×60

=240ml/he

12. Given

50mg=250ml(50000mcg/250ml)

Infusing at 20ml/hr

When 250ml=50000mcg

Then 20ml=??mcg

Cross multiply and divide

=20×50000÷250

=4000mcg/hr

So for 1 minute divide by 60

=66.6 or 67mcg/min

13.Order 50mcg/min

Available 50mg=250ml NS (5000mcg/250ml)

So 50mcg=??ml

Cross multiply and divide

=50×250÷50000

=0.25ml/min

So for 1 hour multiply by 60

=60×0.25

=15ml/hr